OpenStudy (anonymous):

Are my answers right or wrong? Given f(x) = 3x - 1 and g(x)= -x + 6, find f(-2) + g(5). -6 Simplify (3a2)3. 27a^6 Evaluate 60 + 61 + 6² 18 {8}C {7, 8, 9} False

OpenStudy (anonymous):

The first two look right but the evaluation should be a lot bigger, and not sure what the last one is asking

OpenStudy (raden):

the last one is right, {8} is one of the complement of {7,8,9}

OpenStudy (raden):

yea, the 3rd obvious very incorrect. hehe :)

OpenStudy (anonymous):

Oh...thnx

OpenStudy (raden):

is that , your question : 

OpenStudy (anonymous):

Yah

OpenStudy (anonymous):

6^0+6^1+6^2 actually

OpenStudy (raden):

0, 1, and 2as the power right ? oke, 6^0 = 1 6^1 = 6 6^2 = 6 x 6 = 36 so, the total = 1 + 6 + 36 = 43

OpenStudy (anonymous):

Anything to 0 power is 1 so you have 1+6+6^2. 6 Times 6 is 36 so 1+6+36

OpenStudy (raden):

except, 0^0. it is not allowed :)

OpenStudy (anonymous):

That's true haha

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

No problem

OpenStudy (raden):

wlcm